solve the system of equations by the additon method 2x+3y=4
6x+9y=o

Answer by unlockmath(1688)

Hello,
The best way to go about this is multiply the first equation by -3
2x+3y=4
6x+9y=o
Which gives us:
-6x-9y=-12
6x+9y=0
Here is a unique situation because both the x and the y cancel out leaving us when we add both equations:
0 = -12
We know this is impossible so there is no solution. Another way to look at this is to graph both equations. These are straight lines and the two lines do not cross each other. If they crossed each other then exactly where they crossed is the answer for x and y.
Hope this helps.
RJ Toftness
www.math-unlock.com